OpenStudy (anonymous):

Can someone help me with this?? I'm getting confused with the i's. Write a polynomial of the smallest degree with roots 6, 3i and -3i.

OpenStudy (anonymous):

use sum and product for the conjugate roots

OpenStudy (anonymous):

then you will get a quadratic facotr and a linear factor

OpenStudy (anonymous):

Is this the right direction - (x-6) (x-3i) (x-(-3i)

OpenStudy (anonymous):

sum 3i-3i=0 product 3ix-3i=9 quadratic factor = x^2-Sx+P x^2-0x+9 x^2+9 is ur quadratic factor

OpenStudy (anonymous):

(x-6)(x^2+9)

OpenStudy (anonymous):

then expand

OpenStudy (anonymous):

Thanks! I think I got it: x^3-6x^2+3x-54=0